1. One number is 40 more than another. Find the number if the sum is 48.
2. The difference between 6 times a number and 12 equals the sum of the number and thirteen. Find the number.
3. One number exceeds another by 5. If the larger is increased by 5 and the smaller decreased by 8, the larger will be 12 less than twice the smaller. Find the original number.

Answer by checkley77(12844)

1) X-Y=40
X+Y=48 ADD
--------------
2X=88
X=88/2
X=44 ANS.
44-Y=40
Y=44-40
Y=4 ANS.
PROOF:
44+4=48
48=48
------------------
2) 6X-12=X+13
6X-X=13+12
5X=25
X=25/5
X=5 ANS.
PROOF:
6*5-12=5+13
30-12=18
18=18
-----------------------
3) X-Y=5 OR X=Y+5
(X+5)+12=2(Y-8)
Y+5+5+12=2Y-16
Y-2Y=-16-5-5-12
-Y=-38
Y=38 ANS.
X-38=5
X=5+38
X=43 ANS.
PROOF:
43+5+12=2(38-8)
60=2*30
60=60